OpenStudy (anonymous):

x+y=42 80x+64y=2880 How do I graph this!?!?

OpenStudy (johnweldon1993):

Well we CAN put this in slope intercept form (but i advice just use a graphing calculator) So we want to make both equations look like \(\large y = mx + b\) The first is easy \[\large x + y = 42\] We can subtract 'x' from both sides of the equation to get \[\large y = -x + 42\] showing us that the slope would be -1 and the y-intercept would be 42 The second would be \[\large 80x + 64y = 2880\] Same thing...subtract 80x from both sides \[\large 64y = -80x + 2880\] And this time we need to divide everything by 64 because we want 'y' by itself so \[\large y = -\frac{5}{4}x + 45\] So now we have a slope of -5/4 with a y-intercept of 45 So it would look something like 

OpenStudy (anonymous):

where did you get the 42?

OpenStudy (johnweldon1993):

Well, the equations we came up with were \[\large y = -x + 42\] and \[\large y = -\frac{5}{4}x + 45\] The 42 in the first equation as well as the 45 in the second are the y-intercepts of those respective lines...meaning where they cross the 'y-axis' so I labeled them as such

OpenStudy (anonymous):

yeah but when i put it on the graph they dont intersect